# WikiVideo: Article Generation from Multiple Videos

### v1

@@ Abstract
We present the challenging task of automatically creating a high-level Wikipedia-style article that aggregates information from multiple diverse videos about real-world events, such as natural disasters or political elections. Videos are intuitive sources for retrieval-augmented generation (RAG), but most contemporary RAG workflows focus heavily on text while existing methods for video-based summarization focus on low-level scene understanding rather than high-level event semantics. To close this gap, we introduce WikiVideo, a benchmark consisting of expert-written articles and densely annotated videos that provide evidence for articles’ claims, facilitating the integration of video into RAG pipelines and enabling the creation of in-depth content that is grounded in multimodal sources. We further propose Collaborative Article Generation (CAG), a novel interactive method for article creation from multiple videos. CAG leverages an iterative interaction between an r1-style reasoning model and a VideoLLM to draw higher-level inferences about the target event than is possible with VideoLLMs alone, which fixate on low-level visual features. We benchmark state-of-the-art VideoLLMs and CAG in both oracle retrieval and RAG settings and find that CAG consistently outperforms alternative methods, while suggesting intriguing avenues for future work. [^1] Data and code can be found here: https://github.com/alexmartin1722/wikivideo
@@ 1 Introduction
Audiovisual media is becoming an increasingly dominant form of online information consumption. From firsthand, “in the wild” video footage of natural disasters to professionally edited news coverage of major political events, videos serve as rich sources of information for producing factual, grounded articles. Especially for actively unfolding events, grounding articles in video not only can potentially combat misinformation among readers, but can also provide a useful tool for journalists and other writers to quickly synthesize information about new developments. Figure 1 motivates this task, taking an information request and producing a Wikipedia-style article with references to the supporting video content.
@@
In this work, we introduce WikiVideo, a benchmark that builds upon MultiVENT and evaluates the ability to write event-centric articles in the style of a Wikipedia lead (overview) section based only on video content. Given a request about a real-world event, systems must retrieve a set of relevant videos and then generate an article from the videos’ (visual, audio, and OCR) content. WikiVideo consists of 52 events and nearly 400 relevant videos (from a corpus of 109K; avg. 8 relevant/event), with expert-written reference articles that synthesize information about each event across all relevant videos—forcing systems to not only understand high-level information within a single video, but also to synthesize information across multiple videos on the same topic. Systems able to achieve strong results on WikiVideo would be of significant practical use in grounded article generation from multimodal sources, and would enable both the rapid seeding of new Wikipedia articles for actively unfolding events and the enriching of existing articles with audiovisual content.
@@
    We introduce WikiVideo, a new dataset and task for generating articles from multiple videos. WikiVideo is the first benchmark for multi-video article generation, requiring reasoning across audio and visual information, covering 52 events (topics) as depicted in nearly 400 videos that are densely annotated with modality-specific claim grounding annotations, with expert-written reference articles for each event.
@@ 2 Related Work
Video summarization has been studied on small-scale video datasets, such as SumMe (25 videos; Gygli et al., 2014) and VideoSum (50 videos; Song et al., 2015)—considerably smaller than WikiVideo ($\sim 400$ videos). Work on cross-modal summarization that leverages video largely focuses on producing low-level scene descriptions as the summary, since the associated tasks are chiefly concerned with aligning video scenes with caption-like text (He et al., 2023; Lin et al., 2024; Hua et al., 2024). Other work treats video summaries as mere LLM syntheses of frame-level captions (Hua et al., 2024; Zhang et al., 2024b). In contrast, WikiVideo is focused on summaries that provide high-level information supported by video content.
Ren et al. (2025) recently proposed the task of video retrieval augmented generation (videoRAG). While we too explore retrieval in our experiments (section 5), our data and task are considerably different: whereas Ren et al. exclusively use highly polished videos (e.g. documentaries, lectures), much of WikiVideo consists of raw and amateur-edited footage of events in the wild and in real time. Further, whereas they are concerned with short-form question answering, we are concerned with long-form article generation.
Similarly, there is much other video understanding work oriented toward tasks other than summarization, such as retrieval of (Chen & Dolan, 2011; Xu et al., 2016; Anne Hendricks et al., 2017; Wang et al., 2020) question answering about (Jang et al., 2017; Lei et al., 2018; Yu et al., 2019), and recognition of (Zhou et al., 2019; Sanders et al., 2024) low-level video features and concepts that span a few seconds or that exist purely at the frame level.
@@
has largely been studied in text-only settings and as a multi-document summarization task. Early work in this area was performed as part of the DUC and TAC conferences [^2] DUC: https://duc.nist.gov/; TAC: https://tac.nist.gov, including the DUC 2003-2007 multi-document summarization tasks and the TAC 2010 and 2011 Guided Summarization track. Whereas this early work—and many more recent efforts (Hermann et al., 2015; Nallapati et al., 2016; Fabbri et al., 2019; Huang et al., 2024, i.a.)—focused on summarizing news articles, generation of Wikipedia-style articles has received increasing attention. To our knowledge, Sauper & Barzilay (2009) were the first to have attempted this, focusing on generation of full Wikipedia articles by filling learned article templates with sentences extracted from source articles. Similar to us, Liu et al. (2018b) focus on Wikipedia lead sections, taking article titles and a collection of source documents as input and performing abstractive summarization using a decoder-only Transformer and introducing the WikiSum dataset as part of their work. Zhu et al. (2021) also focus on leads, but take a topic modeling-inspired approach, assigning topics to source article paragraphs and conditioning generation of each lead sentence on paragraphs associated with either a single predicted topic or a mixture of topics. Like Sauper & Barzilay (2009), Shao et al. (2024) tackle full Wikipedia article generation, using a complex pipeline that entails (1) surveying related Wikipedia articles, (2) generating perspectival questions and answers via simulated dialogues, (3) using these dialogues to construct an outline for the article, and (4) populating the outline from section titles and headings of source articles retrieved during (2). Concurrently, Yang et al. (2025) explore multimodal article generation, but aimed at incorporating figures into articles rather than synthesizing information from video sources.
@@ 3 WikiVideo Dataset
Figure 2: The WikiVideo curation process. (1) Sentences in Wikipedia lead sections are decomposed into a set of subclaims. (2) Subclaims are grounded in audio, video, and/or OCR evidence. (3) Leads are rewritten to cover all and only the grounded information.
WikiVideo is built using videos from MultiVENT 1.0 (Sanders et al., 2023) and MultiVENT 2.0 (Kriz et al., 2024) that are linked to Wikipedia articles obtained from the January 2025 dump provided by the MegaWika2 dataset (Barham et al., 2025). Our data collection process consists of five steps: (1) initial event and article selection, (2) article claim decomposition, (3) claim correction, (4) claim grounding, and (5) article rewriting. Figure 2 illustrates the core components of the annotation process: decomposition, grounding, and rewriting.
@@
We select an initial set of events from MultiVENT subject to two constraints: the event must (1) have English-language videos associated with it and (2) must have a link to an English Wikipedia article about the event. [^3] As MultiVENT is multilingual, not all events it contains satisfy (1) and many also do not satisfy (2). In total, there are 58 events in MultiVENT satisfying both criteria, supported by 503 associated videos.
For each, we extract the lead section from the linked Wikipedia article to use as the basis for the remainder of our annotation. Lead sections are distinctly suited to our goal of curating high-level articles, as they are explicitly intended to provide a summary of the most important aspects of the entire page. [^4] https://en.wikipedia.org/wiki/Wikipedia:Manual_of_Style/Lead_section
@@
Next, following recent work on claim decomposition (Min et al., 2023; Wanner et al., 2024b; Gunjal & Durrett, 2024, i.a.), we decompose each sentence of the Wikipedia lead section into a set of contextualized, atomic subclaims via few-shot prompting of Qwen 2.5 32B (Qwen et al., 2025). Expert annotators versed in the claim decomposition literature then manually correct these decompositions to ensure atomicity and faithfulness to the original text. [^5] Appendix B has details on the claim decomposition and subclaim correction.
@@
[…]
@@
The final WikiVideo dataset consists of 52 events (topics) spanning 398 videos annotated with grounded subclaims, where each event is associated with a fully grounded, expert-written article. Table 1 provides a summary, with more in Appendix AB.
@@ 4 Article Generation from Multiple Videos
The WikiVideo article generation task takes as input (1) a topic event $T$, (2) a query about $T$, and (3) a set of videos $V=\{v_{1},\ldots,v_{n}\}$ deemed relevant to (i.e. depicting some facet of) $T$. The output is then a natural language article $A_{p}$ generated conditional on $T$, $Q$, and $V$. In this work, we consider two possible sources for $V$: the reference set of videos for $T$ as annotated in MultiVENT 1.0 and 2.0 (the oracle setting) and a set of videos obtained from a retrieval model (the RAG setting).
@@ 4.1 Collaborative Article Generation (CAG)
Conditional text generation from multiple videos faces several challenges that hinder the efficiency and effectiveness of current methods. First, open-source VideoLLMs are generally trained to produce low-level scene descriptions, making extraction of high-level concepts (necessary for complex event understanding) a challenge, even based on a single video—let alone multiple. Second, running inference over multiple long videos is memory-intensive. For instance, in preliminary experiments with several of the VideoLLMs we consider in section 5, even 8 80GB A100s struggled to accommodate a single long video (5+ minutes) at 1 fps, as well as two or more videos at 0.25 fps.
@@
To help address these limitations, we introduce Collaborative Article Generation (CAG; Figure 3), a method for article generation from multiple videos that draws on recent developments in test-time scaling (DeepSeek-AI et al., 2025; Huang et al., 2025; Weller et al., 2025; Jurayj et al., 2025) in addition to the classic notion of relevance feedback from information retrieval (Rocchio, 1971). CAG features three core components: a VideoLLM, a text-based reasoning model, and a text-only LLM.
@@
Next, the reasoning model assists the VideoLLM in producing a refined summary for each video that covers higher-level information about the underlying event. Concretely, the reasoning model is given both $Q$ (here, the name of the target event $T$) and the initial generic summary for a particular video as input, and is then asked to either: (1) return the original summary if the reasoning model deems it to be adequate, or else (2) generate a new prompt seeking additional information about $T$ not attested in the input summary. This new prompt is then used to elicit a refined summary from the VideoLLM—an action we dub RePrompting. The reasoning model can thus be understood as providing a form of relevance feedback on the VideoLLM-generated summary with respect to $Q$ (and thus to $T$). This process is iterative because the reasoning model may in principle RePrompt repeatedly—requesting new summaries from the VideoLLM that are more relevant to $Q$ until it is satisfied with the result. In practice, we enforce a maximum RePrompting iteration budget—analogous to test-time compute budgets for recent reasoning models—and show in section 5 that a higher budget tends to yield higher-quality articles. [^6] Additionally, the reasoning model’s test-time compute budget—the length of its reasoning chains—could in principle be expanded in order to obtain higher-quality prompts, although we leave an exploration of this hyperparameter to future work. Appendix C has further details on VideoLLM and RePrompting prompts.
@@
Once the reasoning model determines that the current query is adequate (or the RePrompting iteration budget is exhausted), the final article is synthesized using a text-only LLM. This model takes as input (1) the original generic summary output by the VideoLLM for each video; (2) all RePrompted VideoLLM queries and their resulting (more event-targeted) summaries; and (3; optionally) an audio transcript of each video. Given these inputs, the model is then instructed to generate the full article about the event $T$. However, we do not provide the explicit topic to the LLM to prevent generating the article based purely on parametric memory. [^7] See for further details Appendix D
We note that in virtue of RePrompting—enabling a reasoning model to iteratively craft prompts for the VideoLLM in order to produce summaries more explicitly targeted to an event of interest ($T$)—CAG goes some way toward mitigating the problem of summaries that are overly focused on low-level descriptions (our first concern). Further, in processing one video at a time, CAG reduces the memory burden of simultaneous processing of multiple long videos (our second concern).
@@ 5 Experiments
We conduct experiments on WikiVideo that (1) compare different VideoLLMs; (2) benchmark CAG against baseline report generation approaches; (3) assess the impact of including raw audio and audio transcripts as input to the text-only LLM; and (4) evaluate the effectiveness of different retrievers in the RAG setting, where $Q$-relevant videos must first be retrieved. (1)-(3) are conducted in the oracle setting, where $V$ consists of all and only the $Q$-relevant videos.
@@
For VideoLLMs, we consider LLaVA-Video-72B (Zhang et al., 2024a), VAST (Chen et al., 2024), InternVideo2.5-8B (Wang et al., 2025), and QwenVL2.5-72B (Bai et al., 2025). We use DeepSeek-R1 distilled to Qwen-32B (DeepSeek-AI et al., 2025) as the reasoning model and Qwen2.5 Qwen et al. (2025) as the text-only LLM that generates the final article.
@@
$\textsc{Concat}_{\textsc{Gen}}$ LLaVA-Video 7.34 1.60 4.78 71.99 19.31 5.08
VAST 16.62 1.71 11.19 80.55 8.04 7.13
@@
$\textsc{Concat}_{\textsc{RePrompt}}$ LLaVA-Video 6.36 1.51 4.22 80.03 21.34 5.50
@@
Table 2: WikiVideo article generation results for CAG and baselines (see §5) without audio inputs (i.e. vision only). CAG obtains the strongest results (bolded) across most metrics.
QwenVL QwenVL+CAG
The video depicts a catastrophic fire engulfing a historic cathedral, likely Notre-Dame in Paris, France. The flames are intense and widespread, with thick smoke billowing into the sky. The iconic spire of the cathedral is visibly on fire, with the structure glowing red from the heat. As the fire rages, the spire begins to weaken and eventually collapses, sending a cascade of debris downward. The surrounding buildings and trees are partially obscured by the smoke, emphasizing the scale and severity of the blaze. The scene is both dramatic and tragic, capturing a moment of significant destruction to a beloved landmark. On April 15, 2019, a catastrophic fire broke out at Notre-Dame Cathedral in Paris, France, causing extensive damage to the historic landmark. The fire, which began in the attic of the cathedral, rapidly spread, engulfing the roof and the iconic spire, which collapsed in a dramatic fall witnessed by a shocked crowd gathered outside. The cathedral, a masterpiece of Gothic architecture dating back to the 12th century, suffered significant structural damage, with the roof and spire completely destroyed. Despite the intense blaze, the main stone structure, including the two bell towers, remained standing. Firefighters worked tirelessly throughout the night to contain the fire, using advanced equipment and techniques to prevent further damage. The event drew global attention, with people around the world expressing sorrow and solidarity. In the aftermath, French President Emmanuel Macron pledged to rebuild the cathedral, and a fundraising campaign was launched to support the restoration efforts. The fire raised questions about the safety and preservation of historic landmarks and highlighted the cultural and spiritual significance of Notre-Dame to both France and the international community.
Table 3: Qualitative Outputs on the Notre-Dame Fire query. The left response comes from video 117864206475218944. Bold text represents information not in the videos.
@@
We use a suite of different metrics to evaluate the generated articles. We first present ROUGE-{1,2,LCS} $\text{F}_{1}$ (R1, R2, RL; Lin, 2004) and BERTScore $\text{F}_{1}$ (BS; Zhang et al., 2019) as two widely used metrics for free-form text generation, using the human-written summaries as references. As these metrics largely focus on lexical similarity, we additionally present AlignScore (AS; Zha et al., 2023), a metric for factual consistency based on a learned text-pair alignment function that outputs a scalar value in $[0,1]$ representing the degree of information alignment between the two texts.
Lastly, we also evaluate the extent to which predicted articles recover specific pieces of event-relevant information. We map each WikiVideo event into the 7-type event ontology defined in the MultiVENT-G dataset (Sanders et al., 2024), each of which is associated with a set of role-like questions about events of that type. [^8] The event types are Sporting Events, Natural Disasters, Elections, Social Events, Demonstrations, Discoveries/Launches, and Political Developments. Event specific scores can be found in Appendix F We use an LLM (GPT-4o) to extract answers to these questions from both the reference and predicted articles. Since a question may have multiple answers, we compute a maximum bipartite matching between predicted and reference answers, obtaining an alignment between them that optimizes normalized edit distance between paired answer spans. We then report an answer span $\text{F}_{1}$ given this alignment, using normalized edit distance in lieu of (overly stringent) exact match. Prior work on event extraction has leveraged similar metrics to evaluate event argument $\text{F}_{1}$ (Du et al., 2021; Chen et al., 2023a; b; Vashishtha et al., 2024), so we refer to this metric as “Arg.”
@@
We consider several baseline article generation methods that ablate different components of CAG. The first baseline ($\textsc{Concat}_{\textsc{Gen}}$) simply concatenates the generic per-video summaries to produce the final article, ablating both the aggregator and reprompting. The second ($\textsc{Concat}_{\textsc{RePrompt}}$) concatenates only the per-video RePrompted summaries, excluding the generic ones, while still ablating the aggregator. The third, (CAG-0), uses the aggregator but fixes CAG’s iteration budget to 0—relying exclusively on the generic per-video summaries. The comparison between CAG-0 and CAG (with an iteration budget of 2) thus offers an illustration of test-time scaling of CAG via a larger iteration budget.
5.1 Experiment 1: CAG and Baselines

### v2

@@ Abstract
We introduce the task of grounded article generation with the goal of creating a Wikipedia-style article from multiple diverse videos about real-world events—from natural disasters to political elections—where all the information in the article is supported by video evidence. Videos are intuitive sources for retrieval-augmented generation (RAG), but most contemporary RAG workflows focus heavily on text while existing methods for video-based summarization focus on low-level scene understanding rather than high-level event semantics. To close this gap, we introduce WikiVideo, a benchmark consisting of expert-written articles and densely annotated videos that provide evidence for articles’ claims, facilitating the integration of video into RAG pipelines and enabling the creation of in-depth content that is grounded in multimodal sources. We further propose Collaborative Article Generation (CAG), a novel interactive method for article creation from multiple videos. CAG leverages an iterative interaction between an r1-style reasoning model and a VideoLLM to draw higher-level inferences about the target event than is possible with VideoLLMs alone, which fixate on low-level visual features. We benchmark state-of-the-art VideoLLMs and CAG in both oracle retrieval and RAG settings and find that CAG consistently outperforms alternative methods, while suggesting intriguing avenues for future work. [^1] Data and code can be found here: https://github.com/alexmartin1722/wikivideo
WikiVideo: Article Generation from Multiple Videos
Alexander Martin1   Reno Kriz1,2*   William Walden1,2*   Kate Sanders1 Hannah Recknor1,2Eugene Yang1Francis Ferraro3Benjamin Van Durme1,2 1Johns Hopkins University 2Human Language Technology Center of Excellence 3University of Maryland Baltimore County {amart233, vandurme}@jhu.edu
@@ 1 Introduction
Audiovisual media is becoming an increasingly dominant form of online information consumption. From firsthand, “in the wild” video footage of natural disasters to professionally edited news coverage of major political events, videos serve as rich sources of information for producing factual, grounded articles. Especially for actively unfolding events, grounding articles in video not only can potentially combat misinformation among readers, but can also provide a useful tool for journalists and other writers to quickly synthesize information about new developments. Figure 1 motivates this task, taking an information request and producing a Wikipedia-style article with references that ground the article in the supporting video content.
@@
In this work, we introduce WikiVideo, a benchmark that builds upon MultiVENT and evaluates the ability to write event-centric articles in the style of a Wikipedia lead (overview) section based only on video content. Given a request about a real-world event, systems must retrieve a set of relevant videos and then generate an article from the videos’ (visual, audio, and OCR) content, providing grounding references (citations) to where the information comes from. WikiVideo consists of 57 events and 427 relevant videos (from a corpus of 109K; avg. 8 relevant/event), with expert-written reference articles that synthesize information about each event across all relevant videos—forcing systems to not only understand high-level information within a single video, but also to synthesize information across multiple videos on the same topic. Systems able to achieve strong results on WikiVideo would be of significant practical use in grounded article generation from multimodal sources, and would enable both the rapid seeding of new Wikipedia articles for actively unfolding events and the enriching of existing articles with audiovisual content.
@@
    We introduce WikiVideo, a new dataset and task for generating articles from multiple videos. WikiVideo is the first benchmark for multi-video article generation, requiring reasoning across audio and visual information, covering 57 events (topics) as depicted in 427 videos that are densely annotated with modality-specific claim grounding annotations, with expert-written reference articles for each event.
@@ 2 Related Work
Video summarization has been studied on small-scale video datasets, such as SumMe (25 videos; Gygli et al., 2014) and VideoSum (50 videos; Song et al., 2015)—considerably smaller than WikiVideo ($\sim 400$ videos). Work on cross-modal summarization that leverages video largely focuses on producing low-level scene descriptions as the summary, since the associated tasks are chiefly concerned with aligning video scenes with caption-like text (He et al., 2023; Lin et al., 2024; Hua et al., 2024). Other work treats video summaries as mere LLM syntheses of frame-level captions (Hua et al., 2024; Zhang et al., 2024a). In contrast, WikiVideo is focused on summaries that provide high-level information supported by video content.
Ren et al. (2025) recently proposed the task of video retrieval augmented generation (videoRAG) over long videos. While we too explore retrieval in our experiments, our data and task are considerably different: whereas Ren et al. exclusively use highly polished videos (e.g. documentaries, lectures) and only single videos, much of WikiVideo consists of raw and amateur-edited footage of events in the wild and in real time and invovles reasoning across information sources. Further, whereas they are concerned with short-form question answering, we are concerned with long-form article generation.
Similarly, there is much other video understanding work oriented toward tasks other than summarization, such as retrieval of (Chen and Dolan, 2011; Xu et al., 2016; Anne Hendricks et al., 2017; Wang et al., 2020) question answering about (Jang et al., 2017; Lei et al., 2018; Yu et al., 2019), and recognition of (Zhou et al., 2019; Sanders et al., 2024) low-level video features and concepts that span a few seconds or exist only at the frame level.
@@
has largely been studied in text-only settings and as a multi-document summarization task. Early work in this area was performed as part of the DUC and TAC conferences [^2] DUC: https://duc.nist.gov/; TAC: https://tac.nist.gov, including the DUC 2003-2007 multi-document summarization tasks and the TAC 2010 and 2011 Guided Summarization track. Whereas this early work—and many more recent efforts (Hermann et al., 2015; Nallapati et al., 2016; Fabbri et al., 2019; Huang et al., 2024, i.a.)—focused on summarizing news articles, generation of Wikipedia-style articles has received increasing attention. To our knowledge, Sauper and Barzilay (2009) were the first to have attempted this, focusing on generation of full Wikipedia articles by filling learned article templates with sentences extracted from source articles. Similar to us, Liu et al. (2018b) focus on Wikipedia lead sections, taking article titles and a collection of source documents as input and performing abstractive summarization using a decoder-only Transformer and introducing the WikiSum dataset as part of their work. Zhu et al. (2021) also focus on leads, but take a topic modeling-inspired approach, assigning topics to source article paragraphs and conditioning generation of each lead sentence on paragraphs associated with either a single predicted topic or a mixture of topics. Like Sauper and Barzilay (2009), Shao et al. (2024) tackle full Wikipedia article generation, using a complex pipeline that entails (1) surveying related Wikipedia articles, (2) generating perspectival questions and answers via simulated dialogues, (3) using these dialogues to construct an outline for the article, and (4) populating the outline from section titles and headings of source articles retrieved during (2). Concurrently, Yang et al. (2025) explore multimodal article generation, but aimed at incorporating figures into articles rather than synthesizing information.
@@
Figure 2: The WikiVideo curation process. (1) Sentences in Wikipedia lead sections are decomposed into subclaims. (2) Subclaims are grounded in audio, video, and/or OCR evidence. (3) Leads are rewritten to cover only the grounded information.
@@ 3 WikiVideo Dataset
WikiVideo is built using videos from MultiVENT 1.0 (Sanders et al., 2023) and MultiVENT 2.0 (Kriz et al., 2024) that are linked to Wikipedia articles obtained from the May 2025 dump provided by the MegaWika2 dataset (Barham et al., 2025). Our data collection process consists of five steps: (1) initial event and article selection, (2) article claim decomposition, (3) claim correction, (4) claim grounding, and (5) article rewriting. Figure 2 illustrates the core components of the annotation process: decomposition, grounding, and rewriting.
@@
We select an initial set of events from MultiVENT subject to two constraints: the event must (1) have English-language videos associated with it and (2) must have a link to an English Wikipedia article about the event. [^3] As MultiVENT is multilingual, not all events it contains satisfy (1) and many also do not satisfy (2). In total, there are 63 events in MultiVENT satisfying both criteria, supported by 503 associated videos.
For each, we extract the lead section from the linked Wikipedia article to use as the basis for the remainder of our annotation. Lead sections are distinctly suited to our goal of curating high-level articles, as they are intended to provide a summary of the most important aspects of the entire page. [^4] https://en.wikipedia.org/wiki/Wikipedia:Manual_of_Style/Lead_section
@@
Next, following recent work on claim decomposition (Min et al., 2023; Wanner et al., 2024b; Gunjal and Durrett, 2024, i.a.), we decompose each sentence of the Wikipedia lead section into a set of contextualized, atomic subclaims via few-shot prompting of Qwen 2.5 32B (Qwen et al., 2025).
Expert annotators versed in the claim decomposition literature then manually correct these decompositions to ensure atomicity and faithfulness to the original text. [^5] Appendix B for details on subclaim decomposition/correction.
@@
Given the corrected subclaims for the Wikipedia lead section associated with each event, we next attempt to ground the subclaims in relevant videos. This grounding task provides modality-specific annotations for each subclaim and for each video associated with the target event, as annotators were asked to indicate whether a subclaim is supported by the video’s (non-text) visual content, its OCR content, its audio content, or none of the above. In-domain expert annotators completed these annotations for all 58 events and 503 videos. Our pilot with these annotators obtained a high overall agreement ($\alpha=.767$).
@@
The final WikiVideo dataset consists of 57 events (topics) spanning 427 videos annotated with grounded subclaims, where each event is associated with a fully grounded, expert-written article. The events in WikiVideo span from 2016-2025, with a specific subset being outside the parametric memory of current (V)LMs. Table 1 provides a summary, with more in Appendix A & B.
@@ 4 Article Generation from Multiple Videos
The WikiVideo article generation task takes as input (1) a topic event $T$, (2) a query about $T$, and (3) a set of videos $V=\{v_{1},\ldots,v_{n}\}$ deemed relevant to (i.e. depicting some facet of) $T$. The output is then a natural language article $A_{p}$ generated conditional on $T$, $Q$, and $V$. $A_{p}$ must be fully grounded in $V$, providing citations to videos that support the claims in $A_{p}$. In this work, we consider two possible sources for $V$: the reference set of videos for $T$ as annotated in MultiVENT 1.0 and 2.0 (the oracle setting) and a set of videos obtained from a retrieval model (the RAG setting).
@@ 4.1 Collaborative Article Generation (CAG)
Conditional text generation from multiple videos faces several challenges that hinder the efficiency and effectiveness of current methods. First, open-source VideoLLMs are generally trained to produce low-level scene descriptions, making extraction of high-level concepts (necessary for complex event understanding) a challenge, even based on a single video—let alone multiple. Second, running inference over multiple long videos is memory-intensive. For instance, in preliminary experiments with several of the VideoLLMs we consider in even 8 80GB A100s struggled to accommodate a single long video (5+ minutes) at 1 fps, as well as two or more videos at 0.25 fps, additionally supported by the findings of Li et al. (2025).
@@
To help address these limitations, we introduce Collaborative Article Generation (CAG; Figure 3), a method for article generation from multiple videos that draws on recent developments in test-time scaling (DeepSeek-AI et al., 2025; Huang et al., 2025; Weller et al., 2025; Jurayj et al., 2025) in addition to the classic notion of relevance feedback from information retrieval (Rocchio, 1971). CAG features three core components: a VideoLLM, a reasoning model, and an LLM.
@@
Next, the reasoning model assists the VideoLLM in producing a refined summary for each video that covers higher-level information about the underlying event. Concretely, the reasoning model is given both $Q$ (here, the name of the target event $T$) and the initial generic summary for a particular video as input, and is then asked to either: (1) return the original summary if the reasoning model deems it to be adequate, or else (2) generate a new prompt seeking additional information about $T$ not attested in the input summary. This new prompt is then used to elicit a refined summary from the VideoLLM—an action we dub RePrompting. The reasoning model can thus be understood as providing a form of relevance feedback on the VideoLLM-generated summary with respect to $Q$ (and thus to $T$). This process is iterative because the reasoning model may in principle RePrompt repeatedly—requesting new summaries from the VideoLLM that are more relevant to $Q$ until it is satisfied with the result. In practice, we enforce a maximum RePrompting iteration budget—analogous to test-time compute budgets for recent reasoning models—that a higher budget tends to yield higher-quality articles.
@@
Once the reasoning model determines that the current query is adequate (or the RePrompting iteration budget is exhausted), the final article is synthesized using a text-only LLM. This model takes as input (1) the original generic summary output by the VideoLLM for each video; (2) all RePrompted VideoLLM queries and their resulting (more event-targeted) summaries; and (3; optionally) an audio transcript of each video. Given these inputs, the model is then instructed to generate the article about $T$.
We note that in virtue of RePrompting—enabling a reasoning model to iteratively craft prompts for the VideoLLM in order to produce summaries more explicitly targeted to an event of interest (TT)—CAG goes some way toward mitigating the problem of summaries that are overly focused on low-level descriptions. Further, in processing one video at a time, CAG reduces the memory burden of simultaneous multi-video inference.
@@ 5 Experiments
We conduct experiments on WikiVideo that (1) benchmark CAG against baseline VLM approaches; (2) demonstrate the strengths of CAG over LLM approaches in generalization to unseen events and grounding information in videos; (3) asses the impact of including raw audio and audio transcripts as input to the text-only LLM; and (4) evaluate the effectiveness of different retrievers in the RAG setting, where videos must first be retrieved. (1-3) are conducted in the oracle setting, where $V$ consists of all and only relevant videos.
@@
For VideoLLMs, we consider LLaVA-Video-72B (“LLaVA-Video” or “LV” in results; Zhang et al., 2024b), VAST (Chen et al., 2024), InternVideo2.5-8B (InternVideo2.5 or IV; Wang et al., 2025), and QwenVL2.5-72B (Qwen2.5VL or QVL; Bai et al., 2025). We use DeepSeek-R1 distilled to Qwen-32B (DeepSeek-AI et al., 2025) as the reasoning model and Qwen2.5 Qwen et al. (2025) as the text-only LLM.
@@
Concat-0 LLaVA-Video 7.34 1.60 4.78 71.99 19.31 5.08
@@
Concat-R LLaVA-Video 6.36 1.51 4.22 80.03 21.34 5.50
@@
Table 2: WikiVideo article generation results for CAG and baselines without audio inputs (i.e. vision only). CAG obtains the best results (bolded) across metrics, with performance increasing with # iterations (2) for most metrics.
@@
We use a suite of different metrics to evaluate the generated articles. We first present ROUGE-{1,2,LCS} $\text{F}_{1}$ (R1, R2, RL; Lin, 2004) and BERTScore $\text{F}_{1}$ (BS; Zhang et al., 2019) as two widely used metrics for free-form text generation, using the human-written articles as references. As these metrics largely focus on lexical similarity, we additionally present AlignScore (AS; Zha et al., 2023), a metric for factual consistency based on a learned text-pair alignment function that outputs a scalar value in $[0,1]$ representing the degree of information alignment between the two texts.
Lastly, we also evaluate the extent to which predicted articles recover specific pieces of event-relevant information. We map each WikiVideo event into the seven-type event ontology defined in the MultiVENT-G dataset (Sanders et al., 2024), each of which is associated with a set of role-focused questions about events of that type. [^6] The event types are Sporting Events, Natural Disasters, Elections, Social Events, Demonstrations, Discoveries/Launches, and Political Developments. Event specific scores in Appendix G. We use an LLM (GPT-4o) to extract answers to these questions from both the reference and predicted articles. Since a question may have multiple answers, we compute a maximum bipartite matching between predicted and reference answers, obtaining an alignment between them that optimizes normalized edit distance between paired answer spans. We then report an answer span $\text{F}_{1}$ given this alignment, using normalized edit distance in lieu of (overly stringent) exact match. Prior work on event extraction has leveraged similar metrics to evaluate event argument $\text{F}_{1}$ (Du et al., 2021; Chen et al., 2023a, b; Vashishtha et al., 2024), so we refer to this as “Arg.”
@@
We consider several baseline article generation methods that ablate different components of CAG. The first baseline (Concat-0) simply concatenates the generic per-video summaries to produce the final article, ablating both the aggregator and reprompting. The second (Concat-RePrompt) concatenates only the per-video RePrompted summaries, excluding the generic ones, while still ablating the aggregator. The third, (CAG-0), uses the aggregator but fixes CAG’s iteration budget to 0—relying exclusively on the generic per-video summaries. The comparison between CAG-0 and CAG-R (iteration budget of 2) thus offers an illustration of test-time scaling of CAG via a larger iteration budget.
5.1 CAG and Baselines